Exercise 2 Silver Holidays markets vacation packages to Los Angeles from Metro Manila. The package includes a round-trip flight on Philippine Airlines (PAL). Silver Holidays pays PAL P60,000 for each round-trip flight. The maximum load on a flight is 300 passengers. Required: 1. What is the unit cost to Silver Holidays of each passenger on a PAL round- trip flight if there are (a) 200, (b)250, or (c) 300 passengers? I play when Silver Holidays is predicting the total air-flight costs to paid next month for PAL carrying 4.000 passengers on 15 scheduled round-trip flights?
Answer #1

1.Amount paid per round-trip flight = P60,000

(a)Number of passengers = 200

Unit Cost per passenger = P60,000/200

= P300

(b) Number of passengers = 250

Unit Cost per passenger = P60,000/250

= P240

(c) Number of passengers = 300

Unit Cost per passenger = P60,000/300

= P200

2.Number of scheduled round-trip flights = 15

Maximum load on a flight = 300

Maximum passengers = 15*300 = 4,500

Passengers to be carried = 4,000

Minimum number of flights needed = 4,000/300 = 14

The unit costs computed in part 1 is of no relevance since 15 flights have already been booked and the charge is fixed per flight.

Total Charges will be 15*60,000 = P900,000

Units costs per passenger will not make any difference to the total air flight costs to be paid.